Production costs chargeable to the Finishing Department in June in Hollins Company are materials $18,260, labor $39,728, overhead $19,100. Equivalent units of production are materials 20,750 and conversion costs 19,675. Production records indicate that 18,600 units were transferred out, and 2,150 units in ending work in process were 50% complete as to conversion cost and 100% complete as to materials.
Cost Reconciliation Cost accounted for Transferred out Work in process Materials Conversion costs Total cost accounted for
Answer #1

Solution:

Cost per equivalent unit

Material =18,260/20,750 = $0.88

Conversion cost =(39,728 +19,100)/19,675 =2.98

Cost reconcilation
Cost accounted for
Transfered out =2.98+0.88=3.86*18,600 71,796
Work in process
Material(2,150*0.88) 1,892
Conversion cost(2,150*0.50*2.98) 3,203 5,095
Total cost accounted for 76,891